# MLIC++: Linear Complexity Multi-Reference Entropy Modeling for Learned Image Compression

### v10

@@ Abstract
The latent representation in learned image compression encompasses channel-wise, local spatial, and global spatial correlations, which are essential for the entropy model to capture for conditional entropy minimization. Efficiently capturing these contexts within a single entropy model, especially in high-resolution image coding, presents a challenge due to the computational complexity of existing global context modules. To address this challenge, we propose the Linear Complexity Multi-Reference Entropy Model (MEM++). Specifically, the latent representation is partitioned into multiple slices. For channel-wise contexts, previously compressed slices serve as the context for compressing a particular slice. For local contexts, we introduce a shifted-window-based checkerboard attention module. This module ensures linear complexity without sacrificing performance. For global contexts, we propose a linear complexity attention mechanism. It captures global correlations by decomposing the softmax operation, enabling the implicit computation of attention maps from previously decoded slices. Using MEM++ as the entropy model, we develop the image compression method MLIC++. Extensive experimental results demonstrate that MLIC++ achieves state-of-the-art performance, reducing BD-rate by $13.39\%$ on the Kodak dataset compared to VTM-17.0 in Peak Signal-to-Noise Ratio (PSNR). Furthermore, MLIC++ exhibits linear computational complexity and memory consumption with resolution, making it highly suitable for high-resolution image coding. Code and pre-trained models are available at https://github.com/JiangWeibeta/MLIC.
@@ 1 Introduction
Recently, various learned image compression models  (Wu et al., 2022; Guo et al., 2022; Ballé et al., 2017; Theis et al., 2017; Ballé et al., 2018; Minnen et al., 2018; Hu et al., 2020; Ma et al., 2020) have emerged, showcasing impressive performance gains. Notably, certain learned image compression models (Minnen & Singh, 2020; Cheng et al., 2020; Zou et al., 2022; Qian et al., 2020; Xie et al., 2021; He et al., 2022; Jiang et al., 2023; Chen et al., 2021; Gao et al., 2021; Chen et al., 2022; Koyuncu et al., 2022; Duan et al., 2023b; Liu et al., 2023; Jiang et al., 2024; Duan et al., 2023a) are already comparable to the advanced traditional method VVC. These models predominantly rely on auto-encoders or variational auto-encoders (Kingma & Welling, 2013), and follow a process that involves transform, quantization, entropy coding, and inverse transform. Entropy coding plays an important role in boosting model performance. An entropy model is utilized to estimate the entropy of the latent representation. A powerful and accurate entropy model usually leads to fewer bits. Expanding contexts of entropy model in learned codecs plays the same role as expanding prediction modes in traditional codecs.
State-of-the-art learned image compression models (Cheng et al., 2020; Minnen & Singh, 2020; Xie et al., 2021; Zhu et al., 2022b; He et al., 2022; Jiang et al., 2023, 2024; Lin et al., 2023) commonly enhance the entropy model by incorporating a hyper-prior module (Ballé et al., 2018) or a context module (Minnen et al., 2018). These additional modules enable the estimation of conditional entropy and the utilization of conditional probabilities for entropy coding. Context modules usually model probabilities and correlations in different dimensions, including local spatial context module, global spatial context module, and channel-wise context module. However, the current global context modules rely on computationally intensive quadratic complexity computations, which consume huge GPU memories and have slower encoding and decoding speed, imposing limitations on the potential for high-resolution image coding. Furthermore, effectively capturing local, global, and channel-wise contexts with acceptable even linear complexity within a single entropy model remains a challenge. To overcome the aforementioned limitations, we propose a novel linear complexity multi-reference entropy model. This entropy model effectively captures local spatial, global spatial, and channel-wise contexts with linear complexity and can be employed for efficient high-resolution image coding, which is denoted as MEM++ to differ from our prior work (Jiang et al., 2023) presented at ACMMM 2023. Based on MEM++, we introduce MLIC++, which achieve state-of-the-art performance.
@@
In comparison to our previous work presented at ACMMM 2023 (Jiang et al., 2023), MLIC++ introduces several significant advancements. The primary distinction lies in the utilization of the proposed linear complexity global spatial context modules without sacrificing performance, as opposed to the quadratic complexity observed in our prior work. This achievement is primarily attributed to the division of the softmax operation, which eliminates the need for a specific order of tensor computation. Our proposed modules incorporate advanced techniques such as learnable position embedding and depth-wise residual bottlenecks (Jiang et al., 2024). Furthermore, MLIC++ captures inter-slice global correlations from all previous slices, in contrast to our previous work (Jiang et al., 2023), which only considers correlations within the previous one slice. MLIC++ exhibits several advantages, including reduced GPU memory consumption and faster encoding and decoding speed. The linear complexity makes our MLIC++ highly suitable for high-resolution image coding. To demonstrate the superiority of MLIC++, we conduct comprehensive performance and complexity comparisons against existing methods (Cheng et al., 2020; Minnen & Singh, 2020; Qian et al., 2020; Xie et al., 2021; Qian et al., 2022; Zhu et al., 2022b; Wang et al., 2022; Zhu et al., 2022a; Zou et al., 2022; He et al., 2022; Liu et al., 2023) across multiple datasets (Toderici et al., 2020; Kodak, 1993; Asuni & Giachetti, 2014) and resolutions (Liu et al., 2020a), extending beyond the preliminary experiments presented in the workshop version (Jiang & Wang, 2023). These advancements in MLIC++ contribute to the field of image compression by offering improved efficiency and performance, while maintaining high-quality compression capabilities.
@@ 3.5 Linear Complexity Intra-Slice Global Context Module
where $\hat{\boldsymbol{y}}^{i-1}_{na,q},\hat{\boldsymbol{y}}^{i-1}_{ac,k}=\textrm{Embedding}\left(\hat{\boldsymbol{y}}^{i-1}\right)$, $\hat{\boldsymbol{y}}^{i}_{ac,v}=\textrm{Embedding}\left(\hat{\boldsymbol{y}}^{i}_{ac}\right)$, Embedding is the embedding layer. Embedding layer consists of a $1\times 1$ convolutional layer and a $3\times 3$ depth-wise convolutional layer. The $3\times 3$ depth-wise convolutional layer is employed for learnable position embedding. This is because the self attention is permutation-invariant and lacks inductive bias. Using a depth-wise convolutional for position embedding does have serval benefits. First, a $3\times 3$ depth-wise convolution is quite light, which has negligible influences on overall complexity. Second, a convolution-based position embedding is flexible for any resolution, due to its translation equivariance. Third, the convolution is able to embed position information because of the zero-padding and the boundary effects (Kayhan & Gemert, 2020; Islam et al., 2019) of images. DepthRB is the depth-wise residual bottleneck (Jiang et al., 2024) and is employed to enhance the non-linearity.
@@
Same as the standard vanilla attention, each row of the implicit similarity matrix ${softmax}_{2}(\hat{\boldsymbol{y}}_{na,q}^{i}){softmax}_{1}(\hat{\boldsymbol{y}}_{na,k}^{i})^{\top}$ sums up to 1 and represents a normalized attention distribution over all positions.
@@ 3.6 Linear Complexity Inter-Slice Global Context Module
where $\hat{\boldsymbol{y}}^{<i}_{q},\hat{\boldsymbol{y}}^{<i}_{k},\hat{\boldsymbol{y}}^{<i}_{v}=\textrm{Embedding}(\hat{\boldsymbol{y}}^{<i})$, Embedding is the embedding layer. Embedding layer consists of a $1\times 1$ convolutional layer and a $3\times 3$ depth-wise convolutional layer. The $3\times 3$ depth-wise convolutional layer is employed for learnable position embedding. DepthRB is the depth-wise residual bottleneck (Jiang et al., 2024) and is employed to enhance the non-linearity. Since the $\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{<i}_{k}\right)^{\top}\hat{\boldsymbol{y}}^{<i}_{v}$ is computed during training and testing, the similarity metric is implicit and the overall complexity is $O(L)$, which is linear with the resolution.
@@ 4.1.1 Training Dataset
Our training datasets contains $10^{5}$ images [^1] https://github.com/JiangWeibeta/MLIC/blob/main/train_list.txt with resolutions exceeding $512\times 512$. These images are selected from ImageNet (Deng et al., 2009), COCO 2017 (Lin et al., 2014), DIV2K (Agustsson & Timofte, 2017), and Flickr2K (Lim et al., 2017). To address the existing compression artifacts in JPEG images, we follow the approach of Ballé et al (Ballé et al., 2018) by further down-sampling the JPEG images using a randomized factor. This downsampling process ensures that the minimum height or width of the images falls within the range of 512 to 584 pixels, effectively reducing the compression artifacts.
@@ 4.2 Benchmarks and Metrics
  • Tecnick (Asuni & Giachetti, 2014) contains 100 $1200\times 1200$ RGB images, which is used as a test set in many methods (Ballé et al., 2018; Minnen et al., 2018; Xie et al., 2021; Kim et al., 2022; Jiang et al., 2023, 2024; Liu et al., 2023; Zhu et al., 2022b; Duan et al., 2023b; Pan et al., 2022).

### v11

@@ Abstract
The latent representation in learned image compression encompasses channel-wise, local spatial, and global spatial correlations, which are essential for the entropy model to capture for conditional entropy minimization. Efficiently capturing these contexts within a single entropy model, especially in high-resolution image coding, presents a challenge due to the computational complexity of existing global context modules. To address this challenge, we propose the Linear Complexity Multi-Reference Entropy Model (MEM++). Specifically, the latent representation is partitioned into multiple slices. For channel-wise contexts, previously compressed slices serve as the context for compressing a particular slice. For local contexts, we introduce a shifted-window-based checkerboard attention module. This module ensures linear complexity without sacrificing performance. For global contexts, we propose a linear complexity attention mechanism. It captures global correlations by decomposing the softmax operation, enabling the implicit computation of attention maps from previously decoded slices. Using MEM++ as the entropy model, we develop the image compression method MLIC++. Extensive experimental results demonstrate that MLIC++ achieves state-of-the-art performance, reducing BD-rate by $13.39\%$ on the Kodak dataset compared to VTM-17.0 in Peak Signal-to-Noise Ratio (PSNR). Furthermore, MLIC++ exhibits linear computational complexity and memory consumption with resolution, making it highly suitable for high-resolution image coding. Code and pre-trained models are available at https://github.com/JiangWeibeta/MLIC. Training dataset is available at https://huggingface.co/datasets/Whiteboat/MLIC-Train-100K.
@@ 1 Introduction
Recently, various learned image compression models  (Wu et al., 2022; Guo et al., 2022; Ballé et al., 2017; Theis et al., 2017; Ballé et al., 2018; Minnen et al., 2018; Hu et al., 2020; Ma et al., 2020) have emerged, showcasing impressive performance gains. Notably, certain learned image compression models (Minnen & Singh, 2020; Cheng et al., 2020; Zou et al., 2022; Qian et al., 2020; Xie et al., 2021; He et al., 2022; Jiang et al., 2023; Chen et al., 2021; Gao et al., 2021; Chen et al., 2022; Koyuncu et al., 2022; Duan et al., 2023b; Liu et al., 2023; Jiang et al., 2024b; Duan et al., 2023a) are already comparable to the advanced traditional method VVC. These models predominantly rely on auto-encoders or variational auto-encoders (Kingma & Welling, 2013), and follow a process that involves transform, quantization, entropy coding, and inverse transform. Entropy coding plays an important role in boosting model performance. An entropy model is utilized to estimate the entropy of the latent representation. A powerful and accurate entropy model usually leads to fewer bits. Expanding contexts of entropy model in learned codecs plays the same role as expanding prediction modes in traditional codecs.
State-of-the-art learned image compression models (Cheng et al., 2020; Minnen & Singh, 2020; Xie et al., 2021; Zhu et al., 2022b; He et al., 2022; Jiang et al., 2023, 2024b; Lin et al., 2023) commonly enhance the entropy model by incorporating a hyper-prior module (Ballé et al., 2018) or a context module (Minnen et al., 2018). These additional modules enable the estimation of conditional entropy and the utilization of conditional probabilities for entropy coding. Context modules usually model probabilities and correlations in different dimensions, including local spatial context module, global spatial context module, and channel-wise context module. However, the current global context modules rely on computationally intensive quadratic complexity computations, which consume huge GPU memories and have slower encoding and decoding speed, imposing limitations on the potential for high-resolution image coding. Furthermore, effectively capturing local, global, and channel-wise contexts with acceptable even linear complexity within a single entropy model remains a challenge. To overcome the aforementioned limitations, we propose a novel linear complexity multi-reference entropy model. This entropy model effectively captures local spatial, global spatial, and channel-wise contexts with linear complexity and can be employed for efficient high-resolution image coding, which is denoted as MEM++ to differ from our prior work (Jiang et al., 2023) presented at ACMMM 2023. Based on MEM++, we introduce MLIC++, which achieve state-of-the-art performance.
@@
In comparison to our previous work presented at ACMMM 2023 (Jiang et al., 2023), MLIC++ introduces several significant advancements. The primary distinction lies in the utilization of the proposed linear complexity global spatial context modules without sacrificing performance, as opposed to the quadratic complexity observed in our prior work. This achievement is primarily attributed to the division of the softmax operation, which eliminates the need for a specific order of tensor computation. Our proposed modules incorporate advanced techniques such as learnable position embedding and depth-wise residual bottlenecks (Jiang et al., 2024b). Furthermore, MLIC++ captures inter-slice global correlations from all previous slices, in contrast to our previous work (Jiang et al., 2023), which only considers correlations within the previous one slice. MLIC++ exhibits several advantages, including reduced GPU memory consumption and faster encoding and decoding speed. The linear complexity makes our MLIC++ highly suitable for high-resolution image coding. To demonstrate the superiority of MLIC++, we conduct comprehensive performance and complexity comparisons against existing methods (Cheng et al., 2020; Minnen & Singh, 2020; Qian et al., 2020; Xie et al., 2021; Qian et al., 2022; Zhu et al., 2022b; Wang et al., 2022; Zhu et al., 2022a; Zou et al., 2022; He et al., 2022; Liu et al., 2023) across multiple datasets (Toderici et al., 2020; Kodak, 1993; Asuni & Giachetti, 2014) and resolutions (Liu et al., 2020a), extending beyond the preliminary experiments presented in the workshop version (Jiang & Wang, 2023). These advancements in MLIC++ contribute to the field of image compression by offering improved efficiency and performance, while maintaining high-quality compression capabilities.
@@ 3.5 Linear Complexity Intra-Slice Global Context Module
where $\hat{\boldsymbol{y}}^{i-1}_{na,q},\hat{\boldsymbol{y}}^{i-1}_{ac,k}=\textrm{Embedding}\left(\hat{\boldsymbol{y}}^{i-1}\right)$, $\hat{\boldsymbol{y}}^{i}_{ac,v}=\textrm{Embedding}\left(\hat{\boldsymbol{y}}^{i}_{ac}\right)$, Embedding is the embedding layer. Embedding layer consists of a $1\times 1$ convolutional layer and a $3\times 3$ depth-wise convolutional layer. The $3\times 3$ depth-wise convolutional layer is employed for learnable position embedding. This is because the self attention is permutation-invariant and lacks inductive bias. Using a depth-wise convolutional for position embedding does have serval benefits. First, a $3\times 3$ depth-wise convolution is quite light, which has negligible influences on overall complexity. Second, a convolution-based position embedding is flexible for any resolution, due to its translation equivariance. Third, the convolution is able to embed position information because of the zero-padding and the boundary effects (Kayhan & Gemert, 2020; Islam et al., 2019) of images. DepthRB is the depth-wise residual bottleneck (Jiang et al., 2024b) and is employed to enhance the non-linearity.
@@
(Jiang et al., 2024a) Same as the standard vanilla attention, each row of the implicit similarity matrix ${softmax}_{2}(\hat{\boldsymbol{y}}_{na,q}^{i}){softmax}_{1}(\hat{\boldsymbol{y}}_{na,k}^{i})^{\top}$ sums up to 1 and represents a normalized attention distribution over all positions.
@@ 3.6 Linear Complexity Inter-Slice Global Context Module
where $\hat{\boldsymbol{y}}^{<i}_{q},\hat{\boldsymbol{y}}^{<i}_{k},\hat{\boldsymbol{y}}^{<i}_{v}=\textrm{Embedding}(\hat{\boldsymbol{y}}^{<i})$, Embedding is the embedding layer. Embedding layer consists of a $1\times 1$ convolutional layer and a $3\times 3$ depth-wise convolutional layer. The $3\times 3$ depth-wise convolutional layer is employed for learnable position embedding. DepthRB is the depth-wise residual bottleneck (Jiang et al., 2024b) and is employed to enhance the non-linearity. Since the $\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{<i}_{k}\right)^{\top}\hat{\boldsymbol{y}}^{<i}_{v}$ is computed during training and testing, the similarity metric is implicit and the overall complexity is $O(L)$, which is linear with the resolution.
@@ 4.1.1 Training Dataset
Our training datasets contains $10^{5}$ images [^1] https://github.com/JiangWeibeta/MLIC/blob/main/train_list.txt [^2] https://huggingface.co/datasets/Whiteboat/MLIC-Train with resolutions exceeding $512\times 512$. These images are selected from ImageNet (Deng et al., 2009), COCO 2017 (Lin et al., 2014), DIV2K (Agustsson & Timofte, 2017), and Flickr2K (Lim et al., 2017). To address the existing compression artifacts in JPEG images, we follow the approach of Ballé et al (Ballé et al., 2018) by further down-sampling the JPEG images using a randomized factor. This downsampling process ensures that the minimum height or width of the images falls within the range of 512 to 584 pixels, effectively reducing the compression artifacts.
@@ 4.2 Benchmarks and Metrics
  • Tecnick (Asuni & Giachetti, 2014) contains 100 $1200\times 1200$ RGB images, which is used as a test set in many methods (Ballé et al., 2018; Minnen et al., 2018; Xie et al., 2021; Kim et al., 2022; Jiang et al., 2023, 2024b; Liu et al., 2023; Zhu et al., 2022b; Duan et al., 2023b; Pan et al., 2022).